OpenStudy (anonymous):

write the following functions as power series, and calculate their open intervals of convergence a) \(\exp(2x^{2})\)

OpenStudy (anonymous):

converges everywhere do you know the expansion of \(e^x\)? substitute \(2x^2\) for \(x\)

OpenStudy (anonymous):

unfortunately i dont know it

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} \frac{x^{n}}{n!}\] is it ?

OpenStudy (anonymous):

yes that is it. need to try to remember this one also sine and cosine

OpenStudy (anonymous):

now it should be easy right? make the substitution

OpenStudy (anonymous):

hmm..

OpenStudy (anonymous):

\[\sum\frac{(2x^2)^n}{n!}\] clean it up a little

OpenStudy (anonymous):

i am sorry i am confused its my second day here, and leraning math since solong :) but math was never so fun before i found this website

OpenStudy (anonymous):

*dont learned math since so long

OpenStudy (anonymous):

do we gonna use \[|\frac{a_{n+1}}{a_{n}}|\] ?

OpenStudy (anonymous):

thats right and R is infinite or for all x is convergence

OpenStudy (anonymous):

aah it was a complete answer?

OpenStudy (anonymous):